500 mL of liquid N2 (density = 0.807 g/mL) is allowed to evaporate into a room where the temperature is 25°C. Assuming a pressure of 1.03 atm, what volume will the gas occupy.
1
Expert's answer
2017-04-13T10:51:04-0400
Using given data, we can define the mass of N2, according to formula:
m (N2) = ρ·V, where ρ is density, V is volume of nitrogen.
Thus:
m (N2) = 0.807·500 = 403.5 (g)
According to Mendeleev – Klapeyron’s equation, we can calculate the volume of gas:
pV = (m(N2))/(M(N2)) ·RT
where p is pressure (p = 1.03 atm); R is gas constant (R = 0,0821 L·atm/mole·K)
Thus, volume of gas equals:
V(N2) = (m(N2))/(M(N2)·p) ·RT = 403.5/(28·1.03) ·0.0821·(273 + 25) = 342.3 (L)

Answer: V(N2) = 342.3 (L).